# Tree-tubings and the combinatorics of resurgent Dyson–Schwinger equations

### v1

@@ 3.1. Rainbows, chains, and other approximations
Certain further restrictions on the class of Feynman diagrams used in the approximation to the perturbative expansion that have been previously considered in the physics literature and specifically in previous analysis of (1) in [5] are the rainbow and chain approximations, the latter of which is important for understanding what are known in physics as renormalons. However, while these two approximations are simpler and so not as hard to understand in and of themselves, it was not entirely clear how to understand the differences between the behaviour of these approximations since they did not live within a common framework.
@@ 3.1.1. Rainbow approximation
These are purely algebraic equations. Let us consider the particular case [^3] For the approximations discussed in Section 3.1, we are particularly interested in the different overall growth behaviour of the coefficients of the expansion and the $m=1$ case is sufficient to illustrate this difference with tidy, relatively easy to understand equations. when $m=1$. In this case the algebraic (albeit nonlinear) equation is $\gamma(x)(1-a_{1}\gamma(x))=x$, with unique solution (scaling to $a_{1}=1$ for convenience)
@@ 3.1.2. Chain approximation
For the chain approximation (as inserted into one outer graph), the insertion trees are corollas, that is trees where every child of the root is a leaf. This is a restriction on the form of the tree being inserted – only single vertex trees can be inserted – so the factor $\gamma(x)$ is replaced by $x$. Additionally, the insertion places must be at the root, so for a corolla with $n$ vertices, we have $n$ insertion places (before all $n-1$ children of the root, between any two adjacent children, and after all children). This results in changing the differential operator from $\left(2x\frac{d}{dx}-1\right)$ to $x\frac{d}{dx}$. All together the system of differential equations for the chain approximation associated to the Mellin transform in (5) is
@@
The right-hand-side is the asymptotic expansion of $-e^{-1/x}\Gamma(0,-1/x)$ as $x\to 0^{+}$, where $\Gamma(a,b)$ is the incomplete gamma function. Notice that unlike the rainbow approximation the expansion coefficients $g_{n}^{\rm chain}$ are factorially divergent. Stirling’s formula for the asymptotic growth rate of $n!$ shows that the exponentially growing prefactor in the asymptotic expansion of $g_{n}^{\rm chain}$ is half the one of the original $a_{1}=1$ Yukawa case in (8). From a resurgence perspective, this means that the associated Borel singularity of $\gamma_{\rm chain}$ is twice as far from the origin as for (8) (see [5]).
@@ 3.1.3. Other intermediate approximations
There are other approximations with natural combinatorial interpretation. For instance, if we change the differential operator from $12x\frac{d}{dx}-$ 1 to $x\frac{d}{dx}$ but keep the $\gamma(x)$ factor, then we have an approximation that is intermediate between the full system and the chain approximation:
@@
The coefficients $g^{\rm odd}_{n}$ are listed in the OEIS as [17, A088716] and they grow factorially.
@@
[…]
Another intermediate approximation is obtained by taking the chain approximation and restoring the differential operator to $12x\frac{d}{dx}-$ 1:
@@
In this case we can interpret the equation as telling us that we insert only single vertices (counted by the leading factor of $x$) but we can insert them into any location in the tree. Again we get all plane rooted trees, but with a restriction on which tubings contribute, namely the only tubings that contribute are those where the lower tube is a single vertex; these are called leaf tubings in [1]. In terms of Feynman diagrams, as with the other intermediate approximation we build the same diagrams as in the full system, but again each diagram’s contribution is somewhat simplified as only some tubings contribute, the specific set of tubings which contribute being different between the two intermediate examples.
@@
The second intermediate approximation has an additional nice property. Since every lower tube is a single vertex in this approximation, all the upper tubes are rooted at the root of the tree. Therefore, the number of compatible labellings of such a tubing of a tree $t$ is simply the number of admissible words of length $|t|-1$. In particular the number of compatible labellings does not depends on the shape of the tree or on the specific tubing (among those tubings under consideration at all in this approximation). Consequently, the coefficient of $x^{n}$ in the anomalous dimension factors into the product of the number of admissible words of length $n-1$ and the number of leaf tubings of rooted trees of size $n$. The first of these factors does not depend on the trees or tubings and the second of these factors does not depend on the Mellin transform of the primitive, so in this example these different contributions separate into different factors at each order.
The number of leaf tubings of a fixed rooted tree is the number of decreasing labellings (or equivalently the number of increasing labellings) of that tree, see Lemma 3.1 of [1]. Increasing labellings of a tree are well-studied objects in combinatorics [2]. However, as we run over all trees, leaf tubings are even easier to count. Given all leaf tubings of all plane rooted trees of size $n$, we can obtain all leaf tubings of all rooted trees of size $n+1$ by inserting a single new leaf in its own tube (and the corresponding new overall tube) in all possible locations in each tubing of each tree. A rooted tree on $n$ vertices has $12n-$ 1 insertion places, so inductively, we obtain that there are $(2n-1)!!$! leaf tubings of plane rooted trees of size $n+1$. This is the same as the number of rooted chord diagrams with $n$ chords.
@@ 3.2. Interpretations for other combinatorial Mellin transforms
With negative rational roots, one can clear denominators as before and only need to consider additionally the presence of factors of the form $(1-b_{i}x)$ in the denominator, with $b_{i}<0$. Taking an alphabet $\Omega_{i}^{\prime}$ with $|b_{i}|$ letters, we can interpret as before, but now there is an overall sign that counts the parity of the number of letters from primed alphabets in the entire compatible labelling. This gives an interpretation, but note that it is blind to any cancellations of terms with opposite sign since this interpretation doesn’t see the signs as signs per se, but merely as another counting variable keeping track of the number of letters from the primed alphabets.
@@
In a slightly different direction, if the Mellin transform is the generating series for some combinatorial class $\mathcal{C}$, then by the same arguments used for the form (5), we can interpret the perturbative solution for the anomalous dimension as counting tubings of plane rooted trees where the tubes with a common root are given the structure of a $\mathcal{C}$-object. This is the same notion of putting a $\mathcal{C}$ structure on something that occurs in the interpretation of composition of generating series [13, Part. A].
@@ 4. Asymptotic resurgent analysis
For applications in both physics and combinatorics, it is interesting to know the asymptotic growth rate of the coefficients $g_{n}$ of the power series $\gamma(x)$ in (4). The previously studied examples [8, 7, 5, 6] showed that the coefficients $g_{n}$ often grow factorially in magnitude and exhibit characteristic resurgent asymptotic behavior. In this section, we prove constraints on this asymptotic behavior directly from the combinatorial tree-tubing result in Theorem 3.5. We do this using a simplified version of alien calculus in the theory of resurgence [12, 15] centered around the ring of factorially divergent power series [3]. Afterwards, we show that the obtained results are consistent with a transseries ansatz, a standard technique to study resurgence behavior [10]. Moroever, we will show that the transseries approach uncovers new behavior if the conditions for the ring of factorially divergent power series are not fulfilled.
@@ 4.1. The ring of factorially divergent power series
    If $f,g\in{\mathbb{R}}^{\alpha}_{\beta}[[x]]$ and $h(x)=f(x)g(x)$, then $h\in{\mathbb{R}}^{\alpha}_{\beta}[[x]]$ and $({\mathcal{A}}^{\alpha}_{\beta}h)(x)=({\mathcal{A}}^{\alpha}_{\beta}f)(x)g(x)+f(x)({\mathcal{A}}^{\alpha}_{\beta}g)(x)$ [3, Proposition 22].
@@
    If $f\in{\mathbb{R}}^{\alpha}_{\beta}[[x]]$, then $\frac{\partial f}{\partial x}\in\frac{1}{x^{2}}{\mathbb{R}}^{\alpha}_{\beta}[[x]]$ and $\left({\mathcal{A}}^{\alpha}_{\beta}\frac{\partial f}{\partial x}\right)(x)=\left(\frac{\alpha^{-1}}{x^{2}}-\frac{\beta}{x}+\frac{\partial}{\partial x}\right)({\mathcal{A}}^{\alpha}_{\beta}f)(x)$ [3, Proposition 15 and 38].
@@
Empirically, we can observe that power series solutions of the differential equation system in Theorem 3.5 are often elements of ${\mathbb{R}}^{\alpha}_{\beta}[[x]]$ for some pair $\alpha,\beta$ [5, 6, 4]. Unfortunately, [3, Proposition 38] does not guarantee that a specific differential equation system has a solution in ${\mathbb{R}}^{\alpha}_{\beta}[[x]]$ (see also the discussion around [3, Proposition 55]), but only provides statements on this solution under the assumption that it is an element of ${\mathbb{R}}^{\alpha}_{\beta}[[x]]$ for some $\alpha,\beta$. Only for the Yukawa Example 3.1 is it relatively straightforward to show rigorously that the power series solution of the differential equation is an element of ${\mathbb{R}}^{\alpha}_{\beta}[[x]]$ for specific values of $\alpha,\beta$ (see [3, Section 7.1]).
@@
(19) $\begin{gathered}({\mathcal{A}}^{\alpha}_{\beta}\gamma_{j})(x)\,-\,({\mathcal{A}}^{\alpha}_{\beta}\gamma_{j-1})(x)\,=\\ a_{j}({\mathcal{A}}^{\alpha}_{\beta}\gamma_{m})(x)\left(2x\frac{\partial}{\partial x}\,-\,1\right)\gamma_{j}(x)\,+\,a_{j}\gamma_{m}(x)\left(2x\frac{\partial}{\partial x}\,-\,1\,+\,2\frac{\alpha^{-1}}{x}\,-\,2\beta\right)({\mathcal{A}}^{\alpha}_{\beta}\gamma_{j})(x)\\ \text{ for }1\leq j\leq m\,.\end{gathered}$
The solution to this system encodes the asymptotic expansions of the power series $\gamma_{j}$ of the form in eq. (14). A simple argument involving the structure of the asymptotic expansion (14) shows that if and only if $f\in{\mathbb{R}}^{\alpha}_{\beta}[[x]]$ and ${\mathcal{A}}^{\alpha}_{\beta}f\in x^{k}{\mathbb{R}}[[x]]$, then $f\in{\mathbb{R}}^{\alpha}_{\beta-k}[[x]]$ and $x^{k}({\mathcal{A}}^{\alpha}_{\beta-k}f)(x)=({\mathcal{A}}^{\alpha}_{\beta}f)(x)$. [3, Proposition 14]. So, assuming that $\gamma_{j}\in{\mathbb{R}}^{\alpha}_{\beta}[[x]]$ for some $\alpha,\beta$ and that the differential system (15) is fulfilled implies that the linear system (19) holds with $({\mathcal{A}}^{\alpha}_{\beta}\gamma_{j})(x)=c_{j}+d_{j}x+\mathcal{O}(x^{2})$ and some specific values for $\alpha,\beta,c_{j}$ and $d_{j}$.
@@
00 0 $\,=\,c_{j}-c_{j-1}-2a_{j}\alpha^{-1}c_{j}$
00 0 $\,=\,d_{j}-d_{j-1}-2a_{j}\alpha^{-1}d_{j}-a_{j}\left(c_{m}-(2\beta+1)c_{j}+2\alpha^{-1}c_{j}\sum_{k=1}^{m}a_{k}\right)\text{ for }1\leq j\leq m\,.$
The set of equations in the first line encodes a linear equation system that only has a nontrivial solution for $c_{j}$ if $\alpha=2a_{k}$ for some $1\leq k\leq m$. From now on for the rest of this section, we require that all $a_{j}$ be distinct and non-vanishing. Hence, for any $1\leq q\leq m$, we have a 1-parameter family of solutions of the first set of equations, $\alpha=2a_{q}$, $c_{k}=0$ for $k<q$, $c_{k}=S_{q}\prod_{j=q+1}^{k}\frac{\alpha}{\alpha-2a_{j}}$ for $q<k\leq m$, and $c_{q}=S_{q}$ an arbitrary value different from 00.
The second set of equations is an inhomogeneous linear equation system in $d_{1},\ldots,d_{m}$. In conjunction with the first set of equations, we may solve directly for $d_{1},\ldots,d_{q-1}$,
@@
(20) 00 0 $=-d_{q-1}-a_{q}\left(c_{m}-(2\beta+1)c_{q}+2\alpha^{-1}c_{q}\sum_{k=1}^{m}a_{k}\right).$
We summarize this discussion of the asymptotic growth rate of the coefficients of the tree-tubing Dyson–Schwinger equations in Theorem 3.5 in the following:
@@
Here, the empty sum is 00, and the empty product is $1$, as usual.
@@
Follows from Theorem 4.1 and the defining Eq. (14) of factorially divergent power series. ∎
@@
An interesting exception is the $m=1$ case, where the prefactor $S$ is known explicitly [3, §7.1].

### v3

@@ 3.1. Rainbows, chains, and other approximations
Certain further restrictions on the class of Feynman diagrams used in the approximation to the perturbative expansion that have been previously considered in the physics literature and specifically in previous analysis of (1) in [8] are the rainbow and chain approximations, the latter of which is important for understanding what are known in physics as renormalons. However, while these two approximations are simpler and so not as hard to understand in and of themselves, it was not entirely clear how to understand the differences between the behavior of these approximations since they did not live within a common framework.
@@ 3.1.1. Rainbow approximation
These are purely algebraic equations. Let us consider the particular case [^3] For the approximations discussed in Section 3.1, we are particularly interested in the different overall growth behavior of the coefficients of the expansion and the $m=1$ case is sufficient to illustrate this difference with tidy, relatively easy to understand equations. when $m=1$. In this case the algebraic (albeit nonlinear) equation is $\gamma(x)(1-a_{1}\gamma(x))=x$, with unique solution (scaling to $a_{1}=1$ for convenience)
@@ 3.1.2. Chain approximation
For the chain approximation (as inserted into one outer graph), the insertion trees are corollas, that is trees where every child of the root is a leaf. This is a restriction on the form of the tree being inserted – only single vertex trees can be inserted – so the factor $\gamma(x)$ is replaced by $x$. Additionally, the insertion places must be at the root, so for a corolla with $n$ vertices, we have $n$ insertion places (before all n1n-1 children of the root, between any two adjacent children, and after all children). This results in changing the differential operator from $\left(2x\frac{d}{dx}-1\right)$ to $x\frac{d}{dx}$. Altogether, the system of differential equations for the chain approximation associated to the Mellin transform in (5) is
@@
The right-hand-side is the asymptotic expansion of $e^{-1/x}\Gamma(0,-1/x)$ as $x\to 0^{+}$, where $\Gamma(a,b)$ is the incomplete gamma function. Notice that unlike the rainbow approximation the expansion coefficients $g_{n}^{\rm chain}$ are factorially divergent. Stirling’s formula for the asymptotic growth rate of $n!$ shows that the exponentially growing prefactor in the asymptotic expansion of $g_{n}^{\rm chain}$ is half the one of the original $a_{1}=1$ Yukawa case in (9). From a resurgence perspective, this means that the associated Borel singularity of $\gamma_{\rm chain}$ is twice as far from the origin as for (9) (see [8]).
@@
The fact that the chain approximation has a differential equation where the appearance of the anomalous dimension on the right hand sides is replaced by the truncation to $x$ was observed in [2] (see equation 7.2 and below) in that case phrased at the level of the Green function and hence giving an equation like the renormalization group equation but with this truncation of the anomalous dimension. This also underlies the special role of chains (there called corollas) in [14].
@@ 3.1.3. Other intermediate approximations
There are other approximations with natural combinatorial interpretation. For instance, if we change the differential operator from $2x\frac{d}{dx}-1$ to $x\frac{d}{dx}$ but keep the $\gamma(x)$ factor, then we have an approximation that is intermediate between the full system and the chain approximation:
@@
The coefficients $g^{\rm odd}_{n}$ are listed in the OEIS as [21, A088716] and they grow factorially. It is interesting to note that this series appears in the study of exponential Dyson-Schwinger equations [1].
@@
with the prefactor $S=0.217950789447151$... which is only known numerically (see, e.g., [9, Appendix A] for a brief explanation on how to obtain such constants empirically). These coefficients have a similar, but faster, factorial rate of growth compared to the chain approximation. As with the chain approximation, the associated Borel singularity is twice as far from the origin as the leading one of the original approximation for the Yukawa case, indicating slower growth by a factor of $2^{-n}$.
Another intermediate approximation is obtained by taking the chain approximation and restoring the differential operator to $2x\frac{d}{dx}-1$:
@@
In this case we can interpret the equation as telling us that we insert only single vertices (counted by the leading factor of $x$) but we can insert them into any location in the tree. Again we get all plane rooted trees, but with a restriction on which tubings contribute, namely the only tubings that contribute are those where each lower tube (excepting the outermost tube) is a single vertex; these are called leaf tubings in [3]. In terms of Feynman diagrams, as with the other intermediate approximation we build the same diagrams as in the full system, but again each diagram’s contribution is somewhat simplified as only some tubings contribute, the specific set of tubings which contribute being different between the two intermediate examples.
@@
The second intermediate approximation has an additional nice property. Since every lower tube is a single vertex in this approximation, all the upper tubes are rooted at the root of the tree. Therefore, the number of compatible labelings of such a tubing of a tree $t$ is simply the number of admissible words of length |t|1|t|-1. In particular the number of compatible labelings does not depend on the shape of the tree or on the specific tubing (among those tubings under consideration at all in this approximation). Consequently, the coefficient of $x^{n}$ in the anomalous dimension factors into the product of the number of admissible words of length n1n-1 and the number of leaf tubings of rooted trees of size $n$. The first of these factors does not depend on the trees or tubings and the second of these factors does not depend on the Mellin transform of the primitive, so in this example these different contributions separate into different factors at each order.
The number of leaf tubings of a fixed rooted tree is the number of decreasing labelings (or equivalently the number of increasing labelings) of that tree, see Lemma 3.1 of [3]. Increasing labelings of a tree are well-studied objects in combinatorics [5]. However, as we run over all trees, leaf tubings are even easier to count. Given all leaf tubings of all plane rooted trees of size $n$, we can obtain all leaf tubings of all rooted trees of size $n+1$ by inserting a single new leaf in its own tube (and the corresponding new overall tube) in all possible locations in each tubing of each tree. A rooted tree on $n$ vertices has 2n12n-1 insertion places, so inductively, we obtain that there are (2n1)!!(2n-1)!! leaf tubings of plane rooted trees of size $n+1$. This is the same as the number of rooted chord diagrams with $n$ chords.
@@ 3.2. Interpretations for other combinatorial Mellin transforms
With negative rational roots, one can clear denominators as before and only need to consider additionally the presence of factors of the form $(1-b_{i}x)$ in the denominator, with $b_{i}<0$. Taking an alphabet $\Omega_{i}^{\prime}$ with $|b_{i}|$ letters, we can interpret as before, but now there is an overall sign that counts the parity of the number of letters from primed alphabets in the entire compatible labeling. This gives an interpretation, but note that it is blind to any cancellations of terms with opposite sign since this interpretation doesn’t see the signs as signs per se, but merely as another counting variable keeping track of the number of letters from the primed alphabets.
@@
In a slightly different direction, if the Mellin transform is the generating series for some combinatorial class $\mathcal{C}$, then by the same arguments used for the form (5), we can interpret the perturbative solution for the anomalous dimension as counting tubings of plane rooted trees where the tubes with a common root are given the structure of a $\mathcal{C}$-object. This is the same notion of putting a $\mathcal{C}$ structure on something that occurs in the interpretation of composition of generating series [17, Part. A].
@@ 4. Asymptotic resurgent analysis
For applications in both physics and combinatorics, it is interesting to know the asymptotic growth rate of the coefficients $g_{n}$ of the power series $\gamma(x)$ in (4). The previously studied examples [11, 10, 8, 9] showed that the coefficients $g_{n}$ often grow factorially in magnitude and exhibit characteristic resurgent asymptotic behavior. In this section, we prove constraints on this asymptotic behavior directly from the combinatorial tree-tubing result in Theorem 3.5. We do this using a simplified version of alien calculus in the theory of resurgence [16, 19] centered around the ring of factorially divergent power series [6]. Afterwards, we show that the obtained results are consistent with a transseries ansatz, a standard technique to study resurgence behavior [12]. Moreover, we will show that the transseries approach uncovers new behavior if the conditions for the ring of factorially divergent power series are not fulfilled.
@@ 4.1. The ring of factorially divergent power series
    ([6, Proposition 22]) If $f,g\in{\mathbb{R}}^{\alpha}_{\beta}[[x]]$ and $h(x)=f(x)g(x)$, then $h\in{\mathbb{R}}^{\alpha}_{\beta}[[x]]$ and
    $({\mathcal{A}}^{\alpha}_{\beta}h)(x)=({\mathcal{A}}^{\alpha}_{\beta}f)(x)g(x)+f(x)({\mathcal{A}}^{\alpha}_{\beta}g)(x).$
@@
    ([6, Proposition 15 and 38]) If $f\in{\mathbb{R}}^{\alpha}_{\beta}[[x]]$, then $\frac{\partial f}{\partial x}\in\frac{1}{x^{2}}{\mathbb{R}}^{\alpha}_{\beta}[[x]]$ and
    $\left({\mathcal{A}}^{\alpha}_{\beta}\frac{\partial f}{\partial x}\right)(x)=\left(\frac{\alpha^{-1}}{x^{2}}-\frac{\beta}{x}+\frac{\partial}{\partial x}\right)({\mathcal{A}}^{\alpha}_{\beta}f)(x).$
@@
Empirically, we can observe that power series solutions of the differential equation system in Theorem 3.5 are often elements of ${\mathbb{R}}^{\alpha}_{\beta}[[x]]$ for some pair $\alpha,\beta$ [8, 9, 7]. Unfortunately, [6, Proposition 38] does not guarantee that a specific differential equation system has a solution in ${\mathbb{R}}^{\alpha}_{\beta}[[x]]$ (see also the discussion around [6, Proposition 55]), but only provides statements on this solution under the assumption that it is an element of ${\mathbb{R}}^{\alpha}_{\beta}[[x]]$ for some $\alpha,\beta$. Only for the Yukawa Example 3.1 is it relatively straightforward to show rigorously that the power series solution of the differential equation is an element of ${\mathbb{R}}^{\alpha}_{\beta}[[x]]$ for specific values of $\alpha,\beta$ (see [6, Section 7.1]).
@@
(21) $\begin{gathered}({\mathcal{A}}^{\alpha}_{\beta}\gamma_{j})(x)\,-\,({\mathcal{A}}^{\alpha}_{\beta}\gamma_{j-1})(x)\,=\\ a_{j}({\mathcal{A}}^{\alpha}_{\beta}\gamma_{m})(x)\left(2x\frac{\partial}{\partial x}\,-\,1\right)\gamma_{j}(x)\\ \,+\,a_{j}\gamma_{m}(x)\left(2x\frac{\partial}{\partial x}\,-\,1\,+\,2\frac{\alpha^{-1}}{x}\,-\,2\beta\right)({\mathcal{A}}^{\alpha}_{\beta}\gamma_{j})(x)\\ \text{ for }1\leq j\leq m\,.\end{gathered}$
The solution to this system encodes the asymptotic expansions of the power series $\gamma_{j}$ of the form in eq. (15). A simple argument involving the structure of the asymptotic expansion (15) shows that if and only if $f\in{\mathbb{R}}^{\alpha}_{\beta}[[x]]$ and ${\mathcal{A}}^{\alpha}_{\beta}f\in x^{k}{\mathbb{R}}[[x]]$, then $f\in{\mathbb{R}}^{\alpha}_{\beta-k}[[x]]$ and $x^{k}({\mathcal{A}}^{\alpha}_{\beta-k}f)(x)=({\mathcal{A}}^{\alpha}_{\beta}f)(x)$. [6, Proposition 14]. So, assuming that $\gamma_{j}\in{\mathbb{R}}^{\alpha}_{\beta}[[x]]$ for some $\alpha,\beta$ and that the differential system (16) is fulfilled implies that the linear system (21) holds with $({\mathcal{A}}^{\alpha}_{\beta}\gamma_{j})(x)=c_{j}+d_{j}x+\mathcal{O}(x^{2})$ and some specific values for $\alpha,\beta,c_{j}$ and $d_{j}$.
@@
$0$ $\,=\,c_{j}-c_{j-1}-2a_{j}\alpha^{-1}c_{j}$
$0$ $\,=\,d_{j}-d_{j-1}-2a_{j}\alpha^{-1}d_{j}-a_{j}\left(c_{m}-(2\beta+1)c_{j}+2\alpha^{-1}c_{j}\sum_{k=1}^{m}a_{k}\right)\text{ for }1\leq j\leq m\,.$
The set of equations in the first line encodes a linear equation system that only has a nontrivial solution for $c_{j}$ if $\alpha=2a_{k}$ for some $1\leq k\leq m$. From now on for the rest of this section, we require that all $a_{j}$ be distinct and non-vanishing. Hence, for any $1\leq q\leq m$, we have a 1-parameter family of solutions of the first set of equations, $\alpha=2a_{q}$, $c_{k}=0$ for $k<q$, $c_{k}=S_{q}\prod_{j=q+1}^{k}\frac{\alpha}{\alpha-2a_{j}}$ for $q<k\leq m$, and $c_{q}=S_{q}$ an arbitrary value different from 0.
The second set is an inhomogeneous linear equation system in $d_{1},\ldots,d_{m}$. In conjunction with the first set of equations, we may solve directly for $d_{1},\ldots,d_{q-1}$,
@@
(22) $0$ $=-d_{q-1}-a_{q}\left(c_{m}-(2\beta+1)c_{q}+2\alpha^{-1}c_{q}\sum_{k=1}^{m}a_{k}\right).$
We summarize this discussion of the asymptotic growth rate of the coefficients of the tree-tubing Dyson–Schwinger equations in Theorem 3.5 in the following theorem.
@@
Here, the empty sum is 0, and the empty product is $1$, as usual.
@@
The statement follows from Theorem 4.1 and the defining Eq. (15) of factorially divergent power series. ∎
@@
An interesting exception is the $m=1$ case, where the prefactor $S$ is known explicitly [6, §7.1].